# LLM-Meta-SR: Learning to Evolve Selection Operators for Symbolic Regression

### v1

@@ 3.5 Incorporating Domain Knowledge into Prompts
Algorithm evolution requires deep domain expertise, which general-purpose LLMs often lack. To compensate, domain knowledge is commonly incorporated into prompts to enhance LLM-based algorithm evolution [52, 35]. For the design of selection operators, we embed the following principles into the prompt to guide the LLM toward generating more effective solutions:
@@ 4 Experimental Settings
Meta-Evolution: The experiments are conducted using GPT-4.1 Mini. For the outer loop, the population size $N$ is set to 20, the number of solutions generated by mutation $M$ is set to 1, and the number of solutions generated by crossover is 19. The total number of generations is set to 20. A length limit of 30 lines is imposed to control operator complexity. For the inner loop, four datasets are selected to evaluate the quality of the generated selection operators. These datasets have OpenML IDs 505, 4544, 588, and 650, corresponding to the four highest-dimensional datasets in the contemporary symbolic regression benchmark (SRBench) [31]. High-dimensional datasets pose greater challenges for symbolic regression algorithms, and optimizing on these is likely to generalize well to both easy and hard problems. The number of generations in the inner loop is set to 30, and the population size is 100. More detailed experimental settings for the inner symbolic regression loop are provided in Appendix D of the supplementary material. All experiments are repeated three times with different random seeds [67] to report the median and confidence intervals.
Symbolic Regression: For symbolic regression, experiments are conducted on 116 out of the 120 datasets in SRBench, excluding the 4 datasets used during the meta-evolution phase to prevent potential data leakage. Each algorithm is evaluated using 10 independent runs per dataset to ensure statistically robust comparisons. The symbolic regression algorithm used is standard genetic programming with linear scaling, which is a widely adopted framework in symbolic regression research [61, 29]. The parameter settings are the same as those used in the inner symbolic regression loop, except that the number of generations is increased to 100.
@@ 5.1 Meta-Evolution Results
In this section, we conduct a series of experiments to evaluate the performance of our proposed meta-evolution framework (LLM-Meta-SR). Specifically, we compare the full method against several ablated variants: without semantic evolution (W/O Semantics), without semantic evolution and bloat control (W/O SE+BC), without domain knowledge (W/O Knowledge), without domain knowledge and semantic evolution (W/O DK+SE), and without domain knowledge, semantic evolution, and bloat control (W/O DK+SE+BC). Since domain knowledge is not always available in practice, we include ablation studies under the setting without domain knowledge to evaluate how the proposed strategy performs in entirely new domains. In the W/O Semantics setting, random selection with identical objective prevention from ReEvo [69] is used as the baseline. In the W/O SE+BC setting, survival selection is replaced with direct population replacement and elitism, following the strategy used in ReEvo [69]. In the W/O Knowledge setting, semantic evolution and bloat control are retained, but the domain knowledge prompt is removed.
@@
Objective Score: The objective score is measured by the average test $R^{2}$ score of each selection operator across four symbolic regression training datasets. For the W/O DK+SE+BC setting, only one run completed within 24 hours and is thus reported. The results in Figure 5, along with the numerical values presented in Table 1, show that the absence of domain knowledge results in the most significant performance drop. This highlights the importance of domain knowledge in guiding effective algorithm evolution, validating the principles introduced in Section 3.5. This performance gain is consistent with the findings of FunSearch [52] and EoH [35], which suggest that LLMs alone often fail to generate sufficiently good solutions without external guidance.
@@
However, domain knowledge alone is insufficient to discover high-performing operators. As shown in Figure 5, semantic evolution plays a critical role in achieving higher objective scores, both with and without domain knowledge. To investigate the role of semantics, we visualize the semantic distribution of solutions over the first five generations in Figure 6. Each point represents a solution, colored by its average score across the four tasks. Positions are computed via t-SNE based on the four-dimensional score vector. The figure reveals that the number of solutions achieving top-3 performance on at least one dataset consistently exceeds three, indicating that top-performing individuals in terms of average score do not dominate across all tasks. Additionally, the final subplot in Figure 6 shows that some solutions, marked with stars, achieve top-3 performance on at least one dataset despite having modest average scores. This illustrates that relying solely on average performance may overlook individuals with strong task-specific capabilities. These findings highlight the importance of semantic-based algorithm evolution in capturing diverse and task-specific behaviors that are obscured by aggregate metrics.
@@
Table 1: Historical best scores and corresponding code lengths for each algorithm.
@@ 5.2 Discovered Operators
In this section, we present experiments on datasets from contemporary symbolic regression benchmarks to evaluate the performance of the discovered operators. Since the LLM-designed operator incorporates several desirable characteristics of selection operators—including diversity-awareness and stage-awareness—it is referred to as Omni selection (Omni). The code for the Omni selection operator is provided in Appendix B of the supplementary material. We compare the performance of the LLM-generated selection operator against several expert-designed baselines, including automatic $\epsilon$-lexicase selection (AutoLex) [30], probabilistic lexicase selection (PLex) [13], lexicase-like selection via diverse aggregation (DALex) [45], $\epsilon$-lexicase selection with dynamic split (D-Split) [25], random sampling with tournament selection (RDS-Tour) [20], complementary phenotype selection (CPS) [14], tournament selection with sizes 3 and 7 [2], and Boltzmann selection with temperature scheduling (Boltzmann) [56]. All of these operators were manually designed by domain experts.
@@
Figure 9: Tree sizes of different selection operators on symbolic regression benchmarks.
Figure 10: Training times of different selection operators on symbolic regression benchmarks.
Model Size: The distribution of model sizes is presented in Figure 10. The results show that the evolved operator produces smaller models compared to top-performing selection operators like AutoLex and CPS. This is primarily because the evolved operator explicitly incorporates model size into the selection process, biasing the search toward regions where symbolic expressions are more compact. As shown in Figure 10 and Figure 8, model size and performance are not always in conflict. By carefully designing the selection operator, it is possible to evolve symbolic models that achieve both high accuracy and small model size, indicating high interpretability.

### v2

@@ 3.5 Incorporating Domain Knowledge into Prompts
Algorithm evolution requires deep domain expertise, which general-purpose LLMs often lack. To compensate, domain knowledge is commonly incorporated into prompts to enhance LLM-based algorithm evolution (Romera-Paredes et al. 2024; Liu et al. 2024a). For the design of selection operators, we embed the following principles into the prompt to guide the LLM toward generating more effective solutions:
@@ 4 Experimental Settings
Meta-Evolution: The experiments are conducted using GPT-4.1 Mini. For the outer loop, the population size $N$ is set to 20, the number of solutions generated by mutation $M$ is set to 1, and the number of solutions generated by crossover is 19. The total number of generations is set to 20. A length limit of 30 lines is imposed to control operator complexity. For the inner loop, four datasets are selected to evaluate the quality of the generated selection operators. These datasets have OpenML IDs 505, 4544, 588, and 650, corresponding to the four highest-dimensional datasets in the contemporary symbolic regression benchmark (SRBench) (La Cava et al. 2021). High-dimensional datasets pose greater challenges for symbolic regression algorithms, and optimizing on these is likely to generalize well to both easy and hard problems. The number of generations in the inner loop is set to 30, and the population size is 100. More detailed experimental settings for the inner symbolic regression loop are provided in Appendix A of the supplementary material. All experiments are repeated three times (Yao et al. 2025) with random seeds set to 0, 1, and 2, to report the median and confidence intervals.
Symbolic Regression: For symbolic regression, experiments are conducted on 116 out of the 120 datasets in SRBench, excluding the 4 datasets used during the meta-evolution phase to prevent potential data leakage. Each algorithm is evaluated using 10 independent runs per dataset to ensure statistically robust comparisons. The symbolic regression algorithm used is standard genetic programming with linear scaling, which is a widely adopted framework in symbolic regression research (Virgolin et al. 2021; Kronberger et al. 2022). The parameter settings are the same as those used in the inner symbolic regression loop, except that the number of generations is increased to 100.
@@ 5.1 Meta-Evolution Results
In this section, we evaluate our meta-evolution framework (LLM-Meta-SR) against several ablated variants, including: without semantic evolution (W/O Semantics), without semantic evolution and bloat control (W/O SE+BC), without domain knowledge (W/O Knowledge), without domain knowledge and semantic evolution (W/O DK+SE), and without domain knowledge, semantic evolution, and bloat control (W/O DK+SE+BC). Since domain knowledge is not always available in practice, we include ablation studies under the setting without domain knowledge to evaluate how the proposed strategy performs in entirely new domains. In the W/O Semantics setting, random selection with identical objective prevention from ReEvo (Ye et al. 2024) is used as the baseline. In the W/O SE+BC setting, survival selection is replaced with direct population replacement and elitism, following the strategy used in ReEvo (Ye et al. 2024). In the W/O Knowledge setting, semantic evolution and bloat control are retained, but the domain knowledge prompt is removed.
@@
Objective Score: The objective score is measured by the average test $R^{2}$ score of each selection operator across the four training datasets. For the W/O DK+SE+BC setting, only one run completed within 24 hours and is thus reported. The results in Figure 5, along with the numerical values presented in Table 1, show that the absence of domain knowledge causes the largest performance drop, highlighting the importance of domain knowledge introduced in Section 3.5 in guiding algorithm evolution. This is consistent with findings from FunSearch (Romera-Paredes et al. 2024) and EoH (Liu et al. 2024a) that LLMs often require external knowledge for guidance.
@@
However, domain knowledge alone is insufficient to discover high-performing operators. As shown in Figure 5, semantic evolution plays a critical role in achieving higher objective scores, both with and without domain knowledge. To investigate the role of semantics, we visualize the semantic distribution of solutions over the first five generations in Figure 6. Each point represents a solution, colored by its average score across the four tasks. Positions are computed via t-SNE based on the four-dimensional score vector. The figure reveals that the number of solutions achieving top-3 performance on at least one dataset consistently exceeds three, indicating that top individuals by average score do not dominate all tasks. Additionally, the final subplot in Figure 6 shows that some solutions, marked with stars, achieve top-3 performance on at least one dataset despite having modest average scores. This illustrates that relying solely on average performance may overlook individuals with strong task-specific capabilities. To further understand how semantic-aware selection works, a real example of parent crossover is provided in Appendix D of the supplementary material, which further confirms the importance of semantic-aware selection for generating better offspring. These findings highlight the importance of semantics-based algorithm evolution in capturing diverse and task-specific behaviors that aggregate metrics may obscure.
@@
Table 1: Median historical best scores and the corresponding code lengths for each algorithm over three runs.
@@ 5.2 Discovered Operators
In this section, we present experiments on datasets from contemporary symbolic regression benchmarks to evaluate the performance of the discovered operators. Since the LLM-designed operator incorporates several desirable characteristics of selection operators—including diversity-awareness and stage-awareness—it is referred to as Omni selection (Omni). The code for the Omni selection operator is provided in Appendix C of the supplementary material. We compare the performance of the LLM-generated selection operator against several expert-designed baselines, including automatic $\epsilon$-lexicase selection (AutoLex) (La Cava et al. 2019), probabilistic lexicase selection (PLex) (Ding, Pantridge, and Spector 2023), lexicase-like selection via diverse aggregation (DALex) (Ni, Ding, and Spector 2024), $\epsilon$-lexicase selection with dynamic split (D-Split) (Imai Aldeia, De França, and La Cava 2024), random sampling with tournament selection (RDS-Tour) (Geiger, Sobania, and Rothlauf 2025), complementary phenotype selection (CPS) (Dolin, Arenas, and Merelo 2002), tournament selection with sizes 3 and 7 (Banzhaf et al. 1998), and Boltzmann selection with temperature scheduling (Boltzmann) (Shojaee et al. 2025). All of these operators were manually designed by domain experts. This section only presents results for a single LLM-discovered operator. Additional results for other LLM-discovered operators are presented in Appendix E of the supplementary material.
@@
Figure 9: Tree sizes of selection operators on symbolic regression benchmarks.
Figure 10: Training times of selection operators on symbolic regression benchmarks.
Figure 11: Median $R^{2}$ scores, model sizes, and training time of 26 algorithms on the symbolic regression benchmark.
Tree Size: The distribution of model sizes is presented in Figure 10. The results show that the evolved operator produces smaller models compared to top-performing selection operators like AutoLex and CPS. This is primarily because the evolved operator incorporates model size into the selection process, biasing the search toward regions where symbolic expressions are more compact. As shown in Figure 10 and Figure 8, model size and performance are not always in conflict. By carefully designing the selection operator, it is possible to evolve symbolic models that achieve both high accuracy and small model size, indicating high interpretability.
@@
5.3 Analysis on State-of-the-Art Symbolic Regression Algorithms
Experimental Settings In this section, we apply the evolved operator to a state-of-the-art Transformer-assisted symbolic regression algorithm, namely retrieval-augmentation-generation-based symbolic regression (RAG-SR) (Zhang et al. 2025), to evaluate the effectiveness of the proposed operator in the context of modern symbolic regression. The only modification is the replacement of the automatic epsilon lexicase selection (La Cava et al. 2019) with the Omni selection operator described in LABEL:lst:omni-selection-repaired in the supplementary material. All other components remain unchanged. The resulting algorithm is referred to as RAG-SR-Omni. For the experimental datasets, following the setup in Section 4, we use 116 out of 120 regression problems, excluding the four used during meta-evolution.
Experimental Results The results are presented in Figure 11. The figure show that RAG-SR-Omni outperforms RAG-SR in terms of median $R^{2}$ scores, model sizes, and training time, achieving the best performance among 26 symbolic regression and machine learning algorithms. These results indicate that the LLM-evolved selection operator can be seamlessly integrated into state-of-the-art symbolic regression methods to further enhance their performance.